OpenStudy (anonymous):

Find the area of the parallelogram with the defining vectors i) a=( 4,-5) and b= (7,2) ii) a= (4,-5,3) and b= (7,2,-4)

OpenStudy (amistre64):

subtract one vector from the other to get it to the origin; is my first thought

OpenStudy (amistre64):

but my first thought aint always too good :)

OpenStudy (amistre64):



OpenStudy (amistre64):

maybe better idea. magnitude of b * height; where height is the "a"sin(a)

OpenStudy (amistre64):

the determinate of the vector matrix of the first one would be good, but not work out to well for the second

OpenStudy (amistre64):

i can recall the dot product stuff for cos(t) but never for sin(t) for some reason

OpenStudy (anonymous):

Thank you!!

OpenStudy (amistre64):

The cross product a × b is defined as a vector c that is perpendicular to both a and b, with a direction given by the right-hand rule and a magnitude equal to the area of the parallelogram that the vectors span. from the wiki i think

OpenStudy (amistre64):

so |axb| seems to be the area of the parallelogram if im reading this right

OpenStudy (amistre64):

now i gotta recall how to cross a 2component vector

OpenStudy (amistre64):

3 components i can do :)

OpenStudy (amistre64):

ii) a= (4 -5 3) and b= (7 2 -4) --------------- x = 20-6 = 14 y=-(-16-21) = 37 z = 8+35 = 43 sqre them add them and sqrt them for magnitude; might wanna use a calcultor :)

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=sqrt%2814%5E2%2B37%5E2%2B43%5E2%29 58.5 thereabouts?

OpenStudy (amistre64):

cross was good; i can wolf the other one for simplicities sake too

OpenStudy (amistre64):

determinant of the vector matrix on 1 is 43, with any luck i recalled that stuff right

OpenStudy (anonymous):

woowww thank you soo soo much!! :)

OpenStudy (amistre64):

do we know if i was right? lol